OpenStudy (anonymous):

What is the correct reason for statement 5? Prove: 3y + 4 + 6(1 + y) = 9y + 10 Statements Reasons 1. 3y + 4 + 6(1 + y) 1. Given 2. 3y + 4 + (6+ 6y) 2. Distributive Property 3. 3y + (4 + 6) + 6y 3. Associative Property 4. 3y + 10 + 6y 4. Addition 5. 3y + 6y + 10 5. 6. 9y + 10 6. Addition A. Multiplication B. Associative Property C. Addition D. Commutative Property

OpenStudy (agent47):

The order of the terms that are being added doesn't matter. What property is that?

OpenStudy (anonymous):

addition

OpenStudy (agent47):

Not exactly. You're not adding yet, you're rearranging the terms.

OpenStudy (texaschic101):

are you looking for the property for step 5 ?

OpenStudy (anonymous):

yes

OpenStudy (agent47):

which property states that a+b=b+a?

OpenStudy (anonymous):

commutative

OpenStudy (texaschic101):

bingo...yes

OpenStudy (agent47):

yes, same goes for your case: 3y + 10 + 6y = 3y + 6y + 10